OpenStudy (anonymous):

(2x-1) + (2x-1) + (x+3) =180 solve for x.

OpenStudy (anonymous):

@jollysailorbold

OpenStudy (anonymous):

so easy just remove all the brackets and you should be able to solve it.

OpenStudy (anonymous):

Im not good at this at all !! haha im dumb !

OpenStudy (anonymous):

2x - 1 + 2x -1 +x + 3 = 180 5x +1 = 180 5x = 179 x = 179/5 final answer use your calculator to find the value and leave in mixed number

OpenStudy (anonymous):

what you mean?

OpenStudy (anonymous):

X = 179 / 5 =...[FIND THE VALUE USING CALCULATOR]

OpenStudy (anonymous):

whats the value?

OpenStudy (anonymous):

I DONT KNOW :) dont be lazy use a CALCULATOR to find out

OpenStudy (anonymous):

i got 35.8

OpenStudy (anonymous):

Leave in fraction form :)

OpenStudy (anonymous):

im confused.

OpenStudy (anonymous):

figured it out (: